### v1

# Adaptive Diffusion Constrained Sampling for Bimanual Robot Manipulation

Abstract
Coordinated multi-arm manipulation requires satisfying multiple simultaneous geometric constraints across high-dimensional configuration spaces, which poses a significant challenge for traditional planning and control methods. In this work, we propose Adaptive Diffusion Constrained Sampling (ADCS), a generative framework that flexibly integrates both equality (e.g., relative and absolute pose constraints) and structured inequality constraints (e.g., proximity to object surfaces) into an energy-based diffusion model. Equality constraints are modeled using dedicated energy networks trained on pose differences in Lie algebra space, while inequality constraints are represented via Signed Distance Functions (SDFs) and encoded into learned constraint embeddings, allowing the model to reason about complex spatial regions. A key innovation of our method is a Transformer-based architecture that learns to weight constraint-specific energy functions at inference time, enabling flexible and context-aware constraint integration. Moreover, we adopt a two-phase sampling strategy that improves precision and sample diversity by combining Langevin dynamics with resampling and density-aware re-weighting. Experimental results on dual-arm manipulation tasks show that ADCS significantly improves sample diversity and generalization across settings demanding precise coordination and adaptive constraint handling.
Keywords: Diffusion Models, Bimanual Manipulation, Constrained Sampling
1 Introduction
Robots operating in human environments must be capable of performing increasingly complex tasks that involve interaction with diverse objects, unstructured scenes, and coordination across multiple end-effectors. As tasks grow in complexity [1], such as moving large furniture, assembling components, or cleaning expansive surfaces, individual manipulators often reach their physical and functional limits. Multi-degree-of-freedom (DoF) systems, such as bimanual robots and mobile manipulators, provide a scalable way to address such limitations by enabling spatial coordination and collaborative manipulation [2].
One core capability that these systems unlock is collaborative transport, where multiple agents work together to move objects that are too large, heavy, or geometrically constrained for a single robot to handle. However, this capability comes at the cost of dramatically increased complexity: planning motion trajectories for such systems requires satisfying a wide range of spatial and physical constraints,e.g., maintaining stable grasps, ensuring collision-free motion among robots and with the environment, and respecting the kinematic limits of each agent. This results in high-dimensional problems with complex nonlinear constraints in configuration and task spaces [3].
Traditional constrained optimization methods and sampling techniques such as Markov Chain Monte Carlo (MCMC) [4] offer principled approaches to constraint satisfaction in such settings. However, their reliance on local gradient information often limits their capacity to explore multi-modal distributions or reason globally about the scene geometry. Recent generative approaches have shown promise; diffusion models were applied to $\mathrm{SE}(3)$ grasp synthesis [5] and to compositional object placement [6], but they typically assume fixed constraint structures and manually specified weighting schemes. Moreover, they often struggle to scale to systems with high-dimensional, interdependent constraints, such as those in multi-arm or mobile manipulation platforms.
To address these limitations, we introduce Adaptive Diffusion Constrained Sampling (ADCS), a generative framework for high-dimensional, constraint-aware sampling in multi-DoF robotic systems. ADCS is designed to integrate both equality and inequality constraints in $\mathrm{SE}(3)$ task space while remaining adaptable to different robot morphologies and task conditions. The key novelty lies in two adaptive mechanisms. First, we introduce an Adaptive Compositional Diffusion, in which the model during training learns to generate feasible $\mathrm{SE}(3)$ poses conditioned on the scene geometry via Signed Distance Fields (SDFs). During sampling, we apply differentiable constraints in joint space using a chain rule, enabling flexible post-hoc adaptation to robot-specific constraints such as joint limits, reachability, and grasp feasibility. Second, we propose a Compositional Weighting Transformer, in which, instead of manually specifying energy-function weights, we incorporate a Transformer-based architecture that learns to dynamically compose task-specific energy terms, allowing the system to prioritize constraints adaptively across varying tasks and contexts, to allow for optimizing challenging tasks with varied co-occurying constraints.
In addition, ADCS uses a sequential sampling strategy that combines Langevin dynamics with density-aware re-weighting, which leads to faster convergence and better precision in constrained sampling. We evaluate ADCS in several collaborative manipulation scenarios, including robot stippling, surface cleaning and object transport, using both simulated and real-world multi-arm systems; multiple Franka Emika Panda arms and a bimanual TIAGo robot. Across these settings, ADCS consistently outperforms baseline approaches in terms of task success, sampling efficiency, generalization to novel scenes, and constraint satisfaction.
To sum up, our main contributions are as follows. (i) We propose ADCS, a diffusion-based framework for constraint-aware generative sampling in multi-DoF robotic systems. (ii) We introduce a Transformer-based mechanism that learns to compose constraints dynamically, eliminating the need for manual tuning of constraint weights. (iii) We formulate a two-stage sampling process that enables the application of task- and robot-specific constraints in joint space, even post training.
2 Preliminaries
Here, we formally define the constrained sampling problem, describe the types of constraints considered, and outline the NLP-based sampling method used to generate training data for our model.
Problem Statement.     We consider a multi-DoF robotic system composed of multiple manipulators operating in a shared workspace. Let $\mathcal{Q}\subset\mathbb{R}^{n\times d}$ denote the joint configuration space, where $n$ denotes the number of robots/end-effectors and $d$ denotes the Dof of each robot, and let $\mathcal{C}=\{c_{i}\}$ be a set of task-specific constraints on configurations $q\in\mathcal{Q}$. Each constraint is either an equality $c_{i}(q)=0$ or an inequality $c_{i}(q)\leq 0$. These constraints capture spatial relationships, kinematic feasibility, and task semantics across the robot and the environment.
Constraint Types.   We consider the following types of constraints: (i) SE(3) Pose Constraints: enforce absolute or relative poses between end-effectors. Given two poses $H_{1},H_{2}\in\mathrm{SE}(3)$, the constraint is $H_{1}^{-1}H_{2}=T_{\text{target}}$; (ii) Orientation Constraints: constrain either a specific axis (e.g., Y-axis of the gripper aligned with Z-axis of the world) or the full rotation $R=R_{\text{goal}}$; (iii) Midpoint Constraints: enforce the midpoint between two end-effectors to match a target (equality) or lie on a surface (inequality); (iv) Signed Distance Constraints: encode proximity to object surfaces for surface contact or collision avoidance; (v) Joint and Self-Collision Constraints: enforce joint limits and prevent inter-link collisions using a set of differentiable bounding spheres.
NLP Sampling.   To generate training data, we use NLP Sampling [4], a method for sampling from the constraint-satisfying region defined by $\mathcal{C}$. Each constraint $c_{i}(q)$ is mapped to a slack term: inequality constraints use $[c_{i}(q)]_{+}$, and equality constraints use $|c_{i}(q)|$, where $[\cdot]_{+}$ denotes the ReLU function. The aggregated slack vector $s(q)$ defines the relaxed energy:
$F_{\gamma,\mu}(q)=\gamma f(q)+\mu\|s(q)\|^{2},$ (1)
where $f(q)$ is a task-specific density and $\gamma,\mu$ balance task likelihood and constraint satisfaction. Sampling proceeds in two stages: first, Gauss-Newton descent minimizes $\|s(q)\|^{2}$ over $K_{\text{down}}$ steps to find a feasible configuration; then, interior sampling—such as manifold-RRT [7] or Langevin dynamics [8]—is applied for $K_{\text{burn}}$ iterations to generate diverse samples from the feasible region. This approach produces high-quality training data for constraint-aware generative modeling.
3 Adaptive Diffusion Constrained Sampling
In this section, we describe our framework for generating diverse, constraint-satisfying solutions in multi-DoF robotic systems. We first introduce our compositional energy-based formulation for modeling constraints in a diffusion framework. We then describe the model architecture, training loss, and how the learned model can be used for efficient sampling at inference time.
3.1 Adaptive Compositional Diffusion
Given a set of constraints $\mathcal{C}=\{c_{i}\}$, we aim to sample robot poses $H\in\mathrm{SE}(3)^{n}$ (for $n$ end-effectors) that satisfy all constraints in $\mathcal{C}$. We model the conditional distribution over poses using an energy-based model (EBM), enabled by the use of diffusion models. Unlike standard DDPMs [8] that generate samples from noise via learned denoisers, we use diffusion training solely to learn a score function for constraint-driven energy modeling based on denoising score matching [9]. For each individual constraint $c\in\mathcal{C}$, we define an energy function $E_{\theta}(H\mid c)$, and the associated probability is given by $p(H\mid c)\propto\exp\bigl{(}-E_{\theta}(H\mid c)\bigr{)}$. We assume these models assign approximately uniform mass over the feasible region of each constraint. To satisfy all constraints jointly, we construct a composed distribution by minimizing the sum of individual energy functions
$H_{0}=\arg\min_{H}\sum_{c\in\mathcal{C}}E_{\theta}(H\mid c).$ (2)
Similarly to Liu et al. [10], the joint diffusion distribution over noisy latent variables $H_{k}$ (at timestep $k$) under all constraints $c_{0},\ldots,c_{N}\in\mathcal{C}$ is given by
$p(H_{k}\mid c_{0},\ldots,c_{n})\propto p(H_{k})\prod_{i=0}^{N}\frac{p(H_{k}\mid c_{i})}{p(H_{k})},$ (3)
which leads to the following form for the composed energy
$E_{\theta}(H_{k}\mid c_{0},\ldots,c_{n})=w_{0}E_{\theta}(H_{k}\mid c_{0})+\sum_{i=1}^{N}w_{i}\left(E_{\theta}(H_{k}\mid c_{0},c_{i})-E_{\theta}(H_{k}\mid c_{0})\right),$ (4)
where $c_{0}$ is a fundamental constraint (e.g., task-space goal), and the remaining energies are conditioned on the satisfaction of $c_{0}$, since in all of our tasks, the end-effector’s relative pose constraints $c_{0}$ should be satisfied. The weights $w_{i}$ determine the influence of each constraint during composition and can be either fixed or learned (see Sec. 3.2).
To enable learning in the Lie group $\mathrm{SE}(3)^{n}$, we inject Gaussian noise in the Lie algebra $\mathfrak{se}(3)^{n}$, to maintain manifold consistency. Specifically, for a clean pose $H$, we apply a perturbation via the exponential map [5]:
$\widehat{H}=H\cdot\text{Expmap}(\epsilon),\quad\epsilon\sim\mathcal{N}(0,\sigma_{k}^{2}I),\quad\epsilon\in\mathbb{R}^{6n},$ (5)
where $\sigma_{k}$ is the noise scale at timestep $k$, and Expmap maps noise vectors in the tangent space to elements of the Lie group. To train the model, we adopt a denoising score matching objective [9], which minimizes the discrepancy between the model’s gradient and the gradient of the Gaussian perturbation distribution. The training loss is
$\mathcal{L}=\frac{1}{L}\sum_{k=0}^{L}\mathbb{E}_{H,\widehat{H}}\left[\left\|\nabla_{\widehat{H}}E_{\theta}(\widehat{H}\mid\mathcal{C},k)-\nabla_{\widehat{H}}\log q(\widehat{H}\mid H,\sigma_{k}^{2}I)\right\|^{2}\right],$ (6)
where $q(\widehat{H}\mid H,\sigma_{k}^{2}I)$ is the Gaussian perturbation distribution in the Lie algebra $\mathfrak{se}(3)^{n}$, and $\nabla_{\widehat{H}}$ denotes the gradient w.r.t. the perturbed pose $\widehat{H}\in\mathrm{SE}(3)^{n}$. Since $\mathrm{SE}(3)^{n}$ is a Lie group, gradients are computed in the associated tangent space via a local reparameterization (details in the Appendix).
This formulation learns an EBM implicitly through a diffusion process. While standard diffusion models learn a generative process via denoising predictions, our framework instead learns the score function, i.e., the gradient of the log-probability—directly. The trained energy function $E_{\theta}(\widehat{H}\mid\mathcal{C},k)$ thus estimates the gradient of the log-density at different noise levels. At inference time, we discard the forward diffusion and use the learned score function as input to Langevin dynamics, iteratively refining a sample by ascending the composed energy landscape as we describe in Sec. 3.3
Figure 1: Overview of the Adaptive Diffusion Constrained Sampling (ADCS) training architecture. The model consists of three components: a constraint feature encoder, an energy network (MLP), and the Compositional Weighting Transformer (CWT). During training, noisy poses in $\mathrm{SE}(3)^{n}$ are used to learn constraint-aware energy functions. At inference, sampling is performed in joint space using Langevin dynamics guided by the learned energy landscape.
3.2 Compositional Weighting Transformer
To enable flexible and context-aware integration of multiple constraint energies, we introduce the Compositional Weighting Transformer (CWT). Rather than statically assigning scalar weights to each constraint model, the CWT learns dynamic composition weights conditioned on the current energy values and their interrelations. The key intuition is that the Transformer [11] enables contextual weighting: it attends to all constraint-specific energy values jointly and can adapt the relative importance of each constraint depending on their magnitudes, types, or combinations. This makes the composition mechanism permutation-invariant (no fixed order of constraints is required) and adaptive to task conditions, as the model can, for instance, downweight a nearly satisfied constraint and prioritize violated ones during inference.
As illustrated in Figure 1, each energy scalar $E_{\theta}(\widehat{H}\mid c_{i})$ corresponding to constraint $c_{i}\in\mathcal{C}$ is treated as a token. These tokens are concatenated and passed through a Transformer, which outputs a set of normalized weights $\{w_{i}\}$. The final composed energy is $E_{\text{total}}(\widehat{H})=\sum_{i}w_{i}E_{\theta}(\widehat{H}\mid c_{i})$. We use positional encodings for the input tokens, but unlike the original Transformer [11], in CWT, those are randomly sampled at each iteration since there is no intrinsic ordering among constraints. The CWT is trained with a fixed maximum number of constraints but can handle arbitrary subsets of active constraints at test time, as it operates over unordered sets of constraint energies. This enables modularity and compositional generalization across tasks and scenes without retraining. Compared to approaches that directly merge constraint features using fully connected layers [12, 13], CWT preserves the identity of each constraint’s energy and defers fusion to the composition stage. This allows individual constraint energies to retain geometric structure and semantics, while enabling adaptive trade-offs between them. CWT is trained jointly with the energy networks and feature encoders using the diffusion-based loss in Eq. 6, without requiring additional supervision. A full training pipeline and implementation details are provided in A.2 and Algorithm 1 in the Appendix.
Figure 2: Visualization of the inference-time sampling process using ALMC and second-order refinement. Starting from Gaussian joint samples , the model iteratively refines configurations using score-based Langevin updates conditioned on constraint energies.
3.3 Sequential Constrained Sampling
At inference time, we use the learned diffusion model to generate joint configurations that satisfy a set of geometric constraints. To this end, we adopt a sequential sampling strategy inspired by Toussaint et al. [4], which combines Annealed Langevin Monte Carlo (ALMC) sampling with resampling and correction steps. The procedure is summarized in Algorithm 2 in the Appendix.
Annealed Langevin Sampling.   In each ALMC call, we first draw a batch of joint configurations $q_{L}^{n_{s}}$ from a standard Gaussian distribution. Using PyTorch Kinematics [14], we compute the corresponding end-effector poses $H_{L}^{n_{s}}\in\mathrm{SE}(3)^{n}$ via forward kinematics. These poses are then passed through the energy network to compute their corresponding energies $E_{\theta}(H_{L}^{n_{s}}\mid\mathcal{C},k)$. Next, we compute the gradient of the energy w.r.t. the joint configuration $q_{k}^{n_{s}}$ using PyTorch’s AutoDiff [15]. This gradient is obtained via the chain rule $\nabla_{q}E_{\theta}=\frac{\partial E_{\theta}(H_{k}\mid\mathcal{C},k)}{\partial H_{k}}\cdot\frac{\partial H_{k}}{\partial q_{k}^{n_{s}}}$. The joint configurations are then updated using a Langevin step. To improve convergence, we adopt a second-order strategy similar to Carvalho et al. [16], but instead of relying solely on first-order gradients, we introduce an approximate Gauss-Newton update during this correction phase. The update direction is computed as $g_{n_{s}}=-(J^{\top}J)^{-1}J^{\top}\nabla_{q}E_{\theta}$, where $J=\frac{\partial H}{\partial q}$ is the Jacobian of the forward kinematics. This approximation accelerates convergence while preserving the structure of the learned data distribution.
Following each ALMC sampling pass, we perform a deterministic second-order refinement step to improve the quality and constraint satisfaction of the sampled configurations. In this phase, we reduce the step size and omit the Gaussian noise, allowing the model to converge more precisely. We apply a Gauss-Newton correction $g_{n_{s}}$ during this step, using approximate second-order information to guide updates in joint space. Importantly, we restrict this correction to the refinement phase to avoid distorting the learned distribution during stochastic sampling, as discussed in [17, 18].
Resampling Phase.   After the initial ALMC sampling and second-order refinement, we perform a resampling step to promote diversity among the generated configurations. We begin by sorting all samples based on their energy values and selecting a fixed number of the top-performing configurations. To avoid over-representing densely clustered solutions, we estimate the sample density $\rho(x)$ using Kernel Density Estimation (KDE) with a Gaussian kernel [19]:
$\rho(x)=\frac{1}{M\,h}\sum_{i=1}^{M}K\left(\frac{x-x_{i}}{h}\right),\quad K(u)=\frac{1}{\sqrt{2\pi}}\exp\left(-\frac{u^{2}}{2}\right),$ (7)
where $M$ is the number of retained samples, $x_{i}$ are the sampled configurations, and $h$ is the kernel bandwidth. Replication weights are assigned inversely proportional to $\rho(x_{i})$: samples in dense regions are duplicated less frequently, while those in sparse regions are upweighted. This reweighting encourages broader exploration of the configuration space, improving the final sample diversity.
4 Experimental Results
We evaluate our method, ADCS, across a suite of constrained motion generation tasks and compare it to representative baselines: Gauss–Newton uses the sampling cost defined in Appendix A.1, with updates computed as $\Delta q=-(J^{\top}J)^{-1}J^{\top}q_{k}$. Diffusion-CCSP [6] follows a similar energy-based formulation, with fixed (uniform) constraint weights and annealed Langevin dynamics as the sampler. NLP-Sampling uses the method from [4]; note that its internal optimization loop is not externally controllable, so we specify only the number of final samples. All methods are evaluated over 10 random seeds (0–9), and we report average metrics across these runs. For fair comparison, Gauss–Newton and CCSP are run for 600 iterations, matching the sampling budget of ADCS.
4.1 Simulation Results
We design six simulation tasks featuring increasingly complex geometric and pose constraints (see Figure 3). All tasks are visualized using the Python package robotic [^1] https://marctoussaint.github.io/robotic/. For each task, we generate 500 samples using each method.
Figure 3: Sampled configurations for Tasks 1–6. Each image shows the robots’ EE poses satisfying various constraints. Red dots indicate end-effector midpoints. Gray regions are constrained surfaces or volumes. Tasks increase in complexity from relative pose and midpoint constraints (Task 1), to obstacle avoidance (Task 2), to full orientation and surface contact via SDFs (Tasks 5–6).
Tasks Summary. All tasks involve a relative pose constraint between the two end-effectors and a midpoint constraint. Tasks 1–3 add increasing complexity. Task 1: relative pose + symmetric midpoint constraint. Task 2: adds a spherical obstacle to avoid. Task 3: adds orientation alignment with the table plane. Tasks 4–6 shift to constrained placement in a region. Task 4: midpoint lies inside a box (defined via bounds), with fixed end-effector orientations. Task 5: replaces the region with a surface defined by a point cloud. Task 6: loosens the orientation constraint to be normal to the surface. Detailed constraint descriptions and metrics are in Appendix A.4.
In Table 3, we report mean, median, and third-quartile values for position and rotation errors across Tasks 1–6. Under a fixed sampling budget (600 iterations), ADCS consistently yields significantly lower constraint satisfaction errors than Gauss–Newton and Diffusion-CCSP. While NLP-Sampling achieves the lowest errors, it serves as an oracle baseline, as it was used to generate the training data for our model. However, it is substantially slower and cannot support tasks with implicit constraints (e.g., those involving SDFs, such as Tasks 5 and 6). In contrast, ADCS offers near-oracle accuracy with significantly improved efficiency and broader applicability.
We also compare the spatial quality of samples generated by NLP Sampling, CCSP, and ADCS in Table 3, evaluating both coverage and uniformity of the resulting distributions. Tasks 1–3 focus on the 3D position of a single end-effector, while Tasks 4–6 assess the midpoint distribution between two arms. Coverage is computed as the fraction of occupied voxels in a predefined 3D grid, while uniformity is measured by the variance in density across voxels—lower variance indicates a more even spread of samples. Due to the relatively coarse voxel grid used (see Appendix A.6 and A.7), absolute coverage values are modest but still reflect comparative performance. As shown in the table, ADCS achieves superior coverage in 5 out of 6 tasks and lower variance in most cases, highlighting its ability to explore constraint-satisfying regions more thoroughly. To assess generalization beyond the training set, we evaluated ADCS on out-of-distribution constraint values. Results in Subsection 4.3 show that our method maintains low constraint satisfaction errors even in these unseen settings, validating its compositional generalization capability.
Relative Position Error (mm) Midpoint Posistion Error (mm) Relative Rotation Error (rad) End-effector Rotaion Error (rad) Time Cost (s)
Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean
Task 1
NLP Sampling* 0.0462 0.0176 0.0668 0.0190 0.0071 0.0267 9.083e-5 4.2642e-5 0.0001 - - - 60.4539
Gauss-Newton 49.947 0.1177 24.944 35.3406 0.0820 24.344 0.0748 0.0001 0.0363 - - - 5.3302
CCSP 23.1474 3.0507 15.2625 20.6335 6.2582 21.8375 0.0264 0.0089 0.0281 - - - 7.8668
ADCS (ours) 0.4746 0.0191 0.0972 0.2697 0.0561 0.1308 0.0008 0.0005 0.0006 - - - 10.1027
Task 2
NLP Sampling* 0.0460 0.0180 0.0660 0.0193 0.0070 0.0292 9.66e-5 4.652e-5 0.0001 - - - 61.4200
Gauss-Newton 106.9 1.8245 122.5 71.531 1.3548 112.12 0.1635 0.0025 0.1654 - - - 6.5329
CCSP 27.3369 5.7985 23.6867 25.4381 9.9820 31.7004 0.0294 0.0128 0.0373 - - - 9.1025
ADCS (ours) 3.1515 0.0221 0.1004 1.9491 0.0547 0.1294 0.0050 0.0003 0.0004 - - - 10.3321
Task 3
NLP Sampling* 0.0455 0.0639 0.0358 0.0230 0.0177 0.0350 9.705e-5 7.012e-5 0.0001 0.0002 0.0002 0.0003 72.8657
Gauss-Newton 136.5462 73.3846 186.7486 84.4612 48.7573 128.8058 0.2012 0.0975 0.2577 0.0507 0.0263 0.0604 7.6832
CCSP 38.9406 12.5211 38.7049 32.3281 11.1668 38.5522 0.0242 0.0101 0.0309 0.0513 0.0341 0.0671 9.3460
ADCS (ours) 4.4175 1.7281 4.8418 1.3967 0.6189 1.6780 0.0058 0.0025 0.0067 0.0027 0.0018 0.0034 10.0325
Task 4
NLP Sampling* 0.0548 0.0176 0.0712 5.0253 1.0019 4.4726 0.00008 0.00003 0.00009 0.00003 0.00001 0.00004 61.3578
Gauss-Newton 216.1300 158.4931 305.5836 65.4829 53.7265 89.2015 0.2889 0.2358 0.0424 0.2297 0.1751 0.3319 7.7406
CCSP 203.0048 170.4515 257.2366 41.7715 21.8248 56.5533 0.0724 0.0458 0.1007 0.1694 0.1334 0.2218 10.3574
ADCS (ours) 2.1288 0.7755 1.9190 0.5845 0.2435 0.6247 0.0021 0.0009 0.0021 0.0013 0.0007 0.0015 11.3247
Task 5
NLP Sampling* - - - - - - - - - - - - -
Gauss-Newton 158.9109 88.9075 229.3474 103.5229 33.1237 140.6337 0.2101 0.1414 0.2861 0.2466 0.1787 0.3529 8.0950
CCSP 83.5345 49.6703 90.2038 40.3970 28.7511 43.9746 0.0721 0.0363 0.0923 0.3238 0.2995 0.4659 13.9051
ADCS (ours) 8.7797 4.1939 5.4900 1.4535 0.5136 0.6962 0.0133 0.0080 0.0111 0.0288 0.0187 0.0213 15.4546
Task 6
NLP Sampling* - - - - - - - - - - - - -
Gauss-Newton 101.1119 27.6524 137.3068 55.0286 4.5290 62.3024 0.1276 0.0470 0.1604 0.1089 0.0734 0.1239 8.0048
CCSP 52.3152 23.7275 58.6257 19.2963 11.3875 23.6884 0.0539 0.0282 0.0691 0.1680 0.0710 0.1411 13.7951
ADCS (ours) 5.1362 1.7652 4.3743 1.6328 0.5571 1.3955 0.0076 0.0037 0.0072 0.0128 0.0110 0.0154 14.9514
Table 1: Comparison of Gauss–Newton, CCSP, and ADCS across Tasks 1–6. For each task and method, we report the relative position error (mm), midpoint position error (mm), relative rotation error (rad), end-effector rotation error (rad), and computation time (s), each summarized by mean, median, and third quartile (Q3). *We consider NLP-Sampling as an oracle.
Data Distribution Coverage Data Distribution Uniformity (Density Variance)
Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6
NLP Sampling 0.0327 0.0321 0.0235 0.0425 - - 0.2330 0.3388 0.5260 126.7000 - -
CCSP 0.0390 0.0310 0.0342 0.41 0.1575 0.2407 0.1224 0.1687 0.1548 13.7969 7.8824 5.0834
ADCS (ours) 0.0452 0.0315 0.0362 0.6 0.595 0.6425 0.1103 0.3669 0.1338 1.9097 2.6821 1.5696
Table 2: Comparison of data uniformity and coverage as sampled by NLP Sampling, CCSP, and ADCS.
Relative Position Error (mm) Midpoint Posistion Error (mm) Relative Rotation Error (rad) End-effector Rotaion Error (rad)
Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean Median Q3
Task 1
ADCS - Fixed Weights 23.1474 3.0507 15.2625 20.6335 6.2582 21.8375 0.0264 0.0089 0.0281 - - -
ADCS - MLP-based Dynamic Weights 22.3517 3.1133 14.7001 20.8284 6.2295 21.0650 0.0275 0.0087 0.0288 - - -
ADCS - CWT 19.8802 1.8213 10.9339 16.7199 3.2276 14.2207 0.0218 0.0062 0.0189 - - -
Task 2
ADCS - Fixed Weights 27.3369 5.7985 23.6867 25.4381 9.9820 31.7004 0.0294 0.0128 0.0373 - - -
ADCS - MLP-based Dynamic Weights 27.6695 5.8037 23.7298 24.9931 10.3324 30.7822 0.0288 0.0130 0.0364 - - -
ADCS - CWT 23.3256 3.4332 17.9259 20.8651 5.6350 21.1443 0.0248 0.0082 0.0270 - - -
Task 3
ADCS - Fixed Weights 40.2740 13.4279 42.5589 34.8348 14.2743 43.4433 0.0318 0.0147 0.0403 0.0515 0.0325 0.0660
ADCS - MLP-based Dynamic Weights 39.6444 13.5843 40.9558 34.6828 14.1637 43.1051 0.0316 0.0149 0.0403 0.0532 0.0329 0.0668
ADCS - CWT 34.9286 10.8732 35.0300 31.9760 10.2701 34.7487 0.0282 0.0120 0.0337 0.0506 0.0370 0.0640
Task 6
ADCS - Fixed Weights 65.1605 28.4774 63.3622 21.9311 13.6046 27.1177 0.0552 0.0299 0.0682 0.1385 0.0633 0.1195
ADCS - MLP-based Dynamic Weights 67.5489 29.2836 73.4915 22.0434 13.1601 27.7264 0.0585 0.0320 0.0724 0.1039 0.0564 0.1014
ADCS - CWT 62.3271 24.2174 62.1286 18.0475 8.9040 20.2202 0.0334 0.0149 0.0393 0.0786 0.0480 0.0857
Table 3: Comparison of constraint composition mechanisms in ADCS: fixed weights, MLP-based dynamic weights, and the proposed CWT. All variants use the same energy and feature networks and exclude task likelihood information during training. CWT consistently yields the best performance across all tasks. Note that there is no sequential sampling involved in this ablation.
4.2 Ablation Studies
We conduct ablations to evaluate the contributions of key components in our method. Specifically, we analyze: (1) the effect of CWT for dynamic constraint composition, (2) the role of KDE for diversity-aware resampling (Appendix A.8), and (3) the impact of sequential sampling strategies (Appendix A.9).
Dynamic Weighting via CWT.   We compare three composition mechanisms for constraint weighting: fixed scalar weights, MLP-based dynamic weights, and our proposed CWT. All models share the same architecture for the energy and feature networks and are trained without task likelihoods to isolate the effect of the weighting mechanism. As shown in Table 3, CWT consistently outperforms the other approaches, achieving lower errors across all tasks, underscoring its ability to adaptively balance constraint satisfaction in diverse compositions.
4.3 Generalization and Robustness Capability
To evaluate the generalization capability of our model, we tested the constraint values of the In-Distribution (ID) and Out-of-Distribution (OOD) data under different tasks. As shown in Figure 4, in Task 1, we set the relative pose to values out of distribution. In Task 3, we tested the different midpoint positions, which were also out of distribution. In Task 4, we both tested the different bounding box sizes (Task 4(a)) and different bounding box types (Task 4(b)), e.g., a circular bounding box. And in Task 5, we altered the object’s position, i.e. the coordinates of the acquired point cloud changed.
Figure 4: Testing results on OOD data across four different tasks.
As shown in Table 4, even when we input data from outside the training set, our model can still produce valid sampling results. For Task 4(a), we reduced the bounding box area to one-fourth. Although the mean error increased, the median and third quartile remained similar. And for Task 4(b), even though we changed the shape of the bounding box, the errors for each constraint did not show a significant increase.
Relative Position Error (mm) Midpoint Posistion Error (mm) Relative Rotation Error (rad) End-effector Rotaion Error (rad)
Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean Median Q3
Task 1
ID 0.4365 0.0133 0.0896 0.2738 0.0430 0.1255 0.0008 0.0004 0.0005 - - -
OOD 0.2178 0.0119 0.0361 0.1293 0.0191 0.0562 0.0006 0.0004 0.0005 - - -
Task 3
ID 4.4175 1.7281 4.8418 1.3967 0.6189 1.6780 0.0058 0.0025 0.0067 0.0027 0.0018 0.0034
OOD 5.5830 1.6920 4.3161 1.6157 0.5937 1.7443 0.0077 0.0026 0.0070 0.0028 0.0020 0.0032
Task 4(a)
ID 2.1288 0.7755 1.9190 0.5845 0.2435 0.6247 0.0021 0.0009 0.0021 0.0013 0.0007 0.0015
OOD 14.3996 0.3898 2.7692 1.0079 0.0022 0.1166 0.0168 0.0006 0.0030 0.0161 0.0004 0.0039
Task 4(b)
OOD 2.8314 0.0254 0.0576 0.3551 0.0011 0.0260 0.0037 0.00008 0.0001 0.0028 0.00002 0.00003
Task 5
ID 4.5119 4.0697 4.6585 0.6092 0.5000 0.6287 0.0078 0.0070 0.0091 0.0186 0.0184 0.0198
OD 4.8049 4.2965 5.0884 0.2737 0.1441 0.2735 0.0072 0.0063 0.0093 0.0187 0.0065 0.0093
Table 4: Sampling results of both ID and OOD datasets to demonstrate the generalization capability.
Figure 5: Point cloud after adding noise.
To assess the robustness of our model, we conducted tests in Task 5 by introducing Gaussian noise to the point cloud data. As illustrated in Figure 5, Gaussian noise with zero mean and a variance of 0.01 was introduced. We evaluated under Gaussian noise with variances of 0.01 and 0.05. As reported in Table 5, the sampling errors did not show a substantial increase in the presence of noise.
Relative Position Error (mm) Midpoint Posistion Error (mm) Relative Rotation Error (rad) End-effector Rotaion Error (rad)
Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean Median Q3
Task 5
Original Point Cloud 4.5119 4.0697 4.6585 0.6092 0.5000 0.6287 0.0078 0.0070 0.0091 0.0186 0.0184 0.0198
Noisy Point Cloud(0.01) 4.0854 2.3550 3.8315 0.5445 0.4563 0.4716 0.0091 0.0104 0.0077 0.0166 0.0148 0.0179
Noisy Point Cloud(0.05) 6.8765 2.7798 6.3495 0.6011 0.4361 0.4714 0.0085 0.0060 0.0087 0.0224 0.0190 0.0259
Table 5: Comparison of sampling results based on original point cloud and noisy point cloud under Task 5.
4.4 Real-World Experiments
We have designed eight different tasks to test ADCS in the real world. Under various specified constraints as defined in A.10, these tasks involve robot stippling [^2] Inspired by https://vhartmann.com/robo-stippling/ using two Franka arms and an object transport task using a TIAGo robot, as shown in Figure 6. We test bimanual robot stippling under various inputs, from given meshes of characters, manually defined regions (circles and rectangles), and raw point clouds from images. We also show that the algorithm is adaptable to different geometric conditions, to infer samples that lie on a slope. The sampling outcome is shown in Figure 7.
Figure 6: The real-world experiment setup involves two Franka robots collaboratively grasping a pen to carry out stippling operations and a TIAGo to carry objects using both hands.
Figure 7: Final stippling layout of 8 different real-world tasks.
5 Related Works
Constrained Sampling.   Sampling under constraints is central to problems in MCMC, optimization, and robotics. With equality constraints, the task reduces to sampling on a manifold [7, 20], while linear inequalities lead to polytope sampling [21]. Toussaint et al. [4] proposed a restart-based NLP sampler for nonlinear constraints in robotics, assuming only local access to cost and constraint functions. In contrast, we focus on sampling from complex, multimodal constrained distributions.
Generative Models for Constrained Optimization.   Deep generative models such as DDPM [8] and SVGD [22] are powerful tools for sampling from multimodal distributions. Their ability to capture disconnected regions makes them attractive for constraint-aware inference [23]. Existing works [24, 6] apply diffusion models to constraint-aware tasks, but often focus on low-dimensional object poses or categorical approximations. Diffusion-based motion planners [16, 25, 26] reframe planning as sampling in energy-based models, but struggle to scale with multiple, nonlinear constraints. Our method addresses this by learning a compositional energy model with transformer-based adaptive weighting for rich constraint integration.
Automatic Weight Tuning.   Designing appropriate constraint weights is tedious and sensitive [5, 27]. Prior methods either optimize weights [28, 29] or learn them from data [30, 31], but these are not directly applicable to diffusion-based models. Recent approaches [5, 6, 16] use fixed weights, whereas our method trains a transformer (CWT) to reason over constraint types and adaptively weight them at inference.
6 Limitation
First, we encode equality/inequality constraints as soft constraints into the energy function, while it shows strong performance compared with learning-based methods, the precision is still inferior to the hard-constrained analytical jacobian method, as in NLP Sampling. Co-optimizing precision and time efficiency could be considered in the future. Second, all constraints in the SE(3) space are currently defined by manually specifying corresponding numerical values, rather than through instructions like ”align both hands”. An interesting direction would be to combine this with Vision-Language Models (VLMs) and define the related constraints through Large Language Models (LLMs). Lastly, the current training process does not incorporate information about the corresponding joint configurations. Future improvements include joint information during the training phase to improve model performance and accelerate convergence during inference.
7 Conclusion
In this paper, we introduced Adaptive Diffusion Constrained Sampling (ADCS), a generative framework designed to effectively integrate equality and inequality geometric constraints within multi-DoF robot manipulation tasks. Our approach uses a dedicated energy network to calculate the cost of equality and inequality constraints, and introduces an SDF network to incorporate external environment perception, while combined with a Transformer-based architecture for dynamically weighting these constraints during inference. Experimental evaluations across six different tasks demonstrated that ADCS outperforms baseline methods, also applicable to scenarios with small feasible areas and strict spatial constraints. Additionally, through ablation studies, we verified that transformer-based dynamic weights, KDE-based resampling, and sequential sampling significantly enhance sample uniformity and performance. Moreover, our sampler can be used to construct a roadmap for enabling motion planning. Finally, our method’s robust generalization was validated on out-of-distribution data, confirming its adaptability and efficiency in realistic robotic applications.
Acknowledgments
This project has received funding from the European Union’s Horizon Europe programme under Grant Agreement No. 101120823, project MANiBOT.
Appendix A Appendix
A.1 Cost Functions for Sampling
We consider these cost functions in the tasks:
(1) SE(3) Cost: Given two points $H_{1}=\begin{bmatrix}R_{1}&t_{1}\\ 0&1\end{bmatrix}$, $H_{2}=\begin{bmatrix}R_{2}&t_{2}\\ 0&1\end{bmatrix}\in SE(3)$, the cost can be defined as $c_{\mathrm{SE3}}=c_{t}+c_{r}$, where $c_{t}=\|t_{1}-t_{2}\|^{2}$, $c_{r}=\text{Log}(R_{1}^{T}R_{2})$ , while Log($\cdot$) maps a rotation matrix $R\in\mathrm{SO}(3)$ into its tangent space Lie‐algebra $\mathfrak{so}(3)$, here we use Theseus [32] to compute this.
(2)SDF Cost: Here we use the SDF value computing by [14] directly, i.e., $c_{\mathrm{sdf}}(H_{t})=\text{SDF}(H_{t})$.
(3) Obstacle Collision Cost: Similar to the work of  Mukadam et al. [33], we instantiate $N$ collision spheres for the robot. By using differentiable PyTorch kinematics [14] to compute the forward kinematics, we obtain the position $p(q,s_{n})$of each collision sphere, where $q$ denotes the robot’s joint configuration. We then compute the distance to the obstacle via a differentiable signed distance function $d(p,obs)$. For any joint configuration, the obstacle collision cost can be expressed as: $c_{obs}(q)=\sum_{i}^{N}(d(p_{i},obs)-\epsilon_{i})$, where $\epsilon$ is the radius of the collision sphere.
(4) Self-Collision Cost: Similar to the obstacle collision cost, we attach a set of collision spheres as a group to each link for collision detection, and then compute the distances between each pair of sphere’s groups as the self-collision cost.
(5) Joint Limit Cost: We constrain the joints within specified ranges by incorporating an L2 norm and introduce a degree $\epsilon$ of redundancy to minimize edge cases as much as possible.
A.2 Training Structure
Training Architecture   In order to make the trained model flexibly applicable to different types of robots, we perform data noising in SE(3) space during the training phase, and then map to the joint space for sampling during the inference phase. In this work, we focus primarily on SE(3) constraint problems, with each constraint energy network implemented as a five-layer multilayer perceptron (MLP). We manually define feature functions $f_{c}(\cdot)$ for SE(3) constraints to capture the target feature. These features are then encoded and input into the energy network. We also incorporate the object’s point-cloud information as an inequality constraint, e.g., ensuring that sampled points lie within a specified region. To enable our model to generalize better across different object shapes, we define a learnable shape code $z_{o}$ as in work [34], and jointly train a network to predict SDF values with given $z_{o}$, replacing a hand-defined feature function, so that model can closely approximate the true SDF values. Here the ground-truth SDF values are computed during training using [14]. Finally, the predicted SDF values together with the shape code $z_{o}$ are fed as inputs into the energy network.
As shown in Alg. 1, our training process consists of four main parts. First, we use the point-cloud encoder $F^{\rm pc}_{\theta}$ to map the raw point cloud $o$ to a latent code $z_{o}$, then feed $\,z_{o}$ and the perturbed data $\hat{H}$ into the SDF network $F^{\rm sdf}_{\theta}$ to predict the SDF values $\hat{s}$, and compute the prediction loss $L_{\rm sdf}(\hat{s},s)$ against the ground-truth SDF $s$. Second, for each predefined geometric or physical constraint $c_{i}$, we extract the corresponding feature $f_{i}$ from $\hat{H}$ and estimate its energy $e_{i}=E_{\theta}(f_{i})$. If a constraint is inactive, we simply set $e_{i}=0$. Third, we concatenate the energies $\{e_{i}\}$ as tokens and feed them into the Transformer $F^{\rm trans}_{\theta}$, which produces a set of weights $\{w_{i}\}$. The total energy is then computed as $E=\sum_{i}e_{i}\,w_{i}$. Finally, we get the $L_{\mathrm{diff}}$ from Eq. 6 and combine the two losses, while back-propagate to update all model parameters $\theta$, thereby improving both SDF accuracy and constraint consistency.
It is worth noting that our model is jointly trained. However, instead of training on all constraints simultaneously in every iteration, we randomly sample a subset or a single constraint for each training step. This strategy allows the model to better capture the relationships and interactions among different constraints.
Given :  $\theta_{0}$: initial params for $F_{\theta}^{\mathrm{sdf}}$,$F_{\theta}^{\mathrm{pc}}$,$F_{\theta}^{\mathrm{feature}}$,$F_{\theta}^{\mathrm{trans}}$, $E_{\theta}$,
$C:\text{constraints set}$,
$N$: maximum number of constraints
1
Datasets :  $\mathcal{D}_{o}:\{o\}$, object’s point cloud,
[…]
2
3for $s\leftarrow 0$ to $S-1$ do
4
5      Sample noise scale $k,\sigma_{k}\leftarrow[0,\dots,L]$;
6       Sample object point cloud $o\in\mathcal{D}_{o}$;
7       Sample $H\sim\mathcal{D}_{H}$;
8       $\epsilon\sim\mathcal{N}(0,\sigma_{k}^{2}I)$;
       $\hat{H}=H\text{ExpMap}(\epsilon)$ ;
        // Add the noise
9       SDF Train
       $z_{o}\leftarrow F_{\theta}^{\mathrm{pc}}(o)$ ;
        // Encode the point cloud
       $\hat{s}\leftarrow F^{\mathrm{sdf}}_{\theta}(z_{o},\,\hat{H})$ ;
        // Predcit the SDF value
       $s\leftarrow\text{SDF}(\hat{H})$ ;
        // Compute the ground-truth SDF value
10       $L_{\mathrm{sdf}}\leftarrow\mathcal{L}(\hat{s},s)$;
11       Energy Net Train
12       for $i=0$ to $N-1$ do
13             if $c_{i}\in C$ then
                   $f_{i}\leftarrow F_{\theta}^{\rm feature}(\hat{H},c_{i})$ ;
                    // Get the features from each constraint
                   $e_{i}\leftarrow E_{\theta}(f_{i};\theta)$ ;
                    // Estimate each energy
14
15             else
                   $e_{i}\leftarrow 0$ ;
                    // Zero padding
16
17
18      Transfomer-based Weights Train
       $\{w_{i}\}_{i=0}^{N-1}\leftarrow F_{\theta}^{\mathrm{trans}}(\{e_{i}\}_{i=0}^{N-1})$ ;
        // Compute the weights using Transformer
19
20      $e_{total}\leftarrow\sum_{i=0}^{N-1}e_{i}\cdot w_{i}$;
       $L_{\mathrm{diff}}=\mathcal{L}(e_{total},\hat{H},H,\sigma_{k})$;
        // Compute the loss $\triangleright$ Eq. 6
21       Parameter Update
22       $L\leftarrow L_{\mathrm{sdf}}+L_{\mathrm{diff}}$;
       $\theta\leftarrow\theta-\eta\,\nabla_{\theta}L$ ;
        // Update the parameters $\theta$
23
Algorithm 1 Adaptive Diffusion Constraint Model Training
A.3 Sampling Structure
Input: $\{\sigma_{k}\}_{k=1}^{L}$: Noise levels;
     $L$: Diffusion steps;
     $\epsilon$: Step rate;
     $n_{s}$: Number of initial samples;
     $n_{r}$: Number of resampling samples;
     $t_{f}$: Fixed time step;
     $\alpha_{f}$: Fixed step rate;
     $C$: Constraint set
Output: Final samples $q_{0}^{n_{s}}$
1 Initialize $q_{L}^{n_{s}}\sim p_{L}(q)$;
$q_{0}^{n_{s}},e_{0}^{n_{s}}\leftarrow\text{ALMC}(q_{L}^{n_{s}},\sigma_{k},L,\epsilon,t_{f},\alpha_{f},C)$ ;
  // Get the joint configurations and energy after fisrt stage sampling $\triangleright$ Alg. 3
$q_{0}^{n_{r}}\leftarrow\text{sort}(q_{0}^{n_{s}},e_{0}^{n_{s}}$) ;
  // Select data with less energy
$\rho(q)\leftarrow\text{KDE}(q_{0}^{n_{r}})$ ;
  // Estimate density
$w_{i}\propto 1/\rho(q_{i})$ ;
  // Compute the replication weights
$q_{L}^{n_{s}}\leftarrow\text{repeat}(q_{0}^{n_{r}},w_{i})$ ;
  // Repeat data based on weights
$q_{0}^{n_{s}},e_{0}^{n_{s}}\leftarrow\text{ALMC}(q_{L}^{n_{s}},\sigma_{k},L,\epsilon,t_{f},\alpha_{f},C)$ ;
  // Get final $q_{0}^{n_{s}}$triangleright Alg. 3
2
Algorithm 2 Adaptive Diffusion Constraint Sampler
Input: $\{\sigma_{k}\}_{k=1}^{L}$, $L$, $\epsilon$, $n_{s}$, $t_{f}$, $\alpha_{f}$, $q_{L}^{n_{s}}$, $C$
Output: Final samples $q_{0}^{n_{s}}$
1
2for $k\leftarrow L$ to $1$ do
3       $H_{k}^{n_{s}}\leftarrow\text{FK}(q_{k}^{n_{s}})$;
       $e_{n_{s}}\leftarrow E_{\theta}(H_{k}^{n_{s}},k,C)$ ;
        // Compute the energy
       $\alpha_{k}\leftarrow\epsilon\cdot\sigma_{k}/\sigma_{L}$ ;
        // Select step size
       $\epsilon\sim\mathcal{N}(0,I)$ ;
        // Sample white noise in joint space
       $q_{k-1}^{n_{s}}\leftarrow q_{k}^{n_{s}}+\left(-\frac{\alpha_{k}^{2}}{2}\frac{\partial e_{n_{s}}}{\partial q_{k}^{n_{s}}}+\alpha_{k}\epsilon\right)$ ;
        // Make a Langevin dynamics step
4
5
6for $k\leftarrow L$ to $1$ do
7       $H_{k}^{n_{s}}\leftarrow\text{FK}(q_{k}^{n_{s}})$;
       $e_{n_{s}}\leftarrow E_{\theta}(H_{k}^{n_{s}},t_{f},C)$ ;
        // Compute the energy
       $e_{g}\leftarrow f(H_{k}^{n_{s}},C)$ ;
        // Compute the sampling cost
       $J_{g}=\frac{\partial e_{g}}{\partial q_{k}^{n_{s}}}$ ;
        // Compute the jacobian
       $g_{n_{s}}\leftarrow(J_{g}^{T}J_{g})^{-1}J^{T}q_{k}^{n_{s}}$ ;
        // Compute the Gauss Newton
       $q_{k-1}^{n_{s}}\leftarrow q_{k}^{n_{s}}+\left(-\frac{\alpha_{f}^{2}}{2}\frac{\partial e_{n_{s}}}{\partial q_{k}^{n_{s}}}-g_{n_{s}}\right)$ ;
        // Make a Langevin dynamics step
8
9
10return $q_{0}^{n_{s}}$;
Algorithm 3 Annealed Langevin Markov Chain Monte Carlo Sampler (ALMC)
Here we describe our sequential sampling structure. It consists of two stages, each of which uses the ALMC algorithm. Between these stages, we perform resampling via a KDE-based replication method. The sampler consists of two nested components:
Adaptive Diffusion Constraint Sampler: Generate $n_{s}$ initial joint configurations from $\mathcal{N}(0,I)$ and run one pass of ALMC to compute their energies at each noise level. Sort these samples by energy, estimate their density via KDE, and compute replication weights to resample $n_{r}$ configurations. Finally, run a second ALMC pass on the resampled set to produce the final samples.
Annealed Langevin MCMC: For each noise level $\sigma_{k}$, perform Langevin updates by computing $\nabla_{q}E$ via the chain rule $\frac{\partial E}{\partial H}\frac{\partial H}{\partial q}$ and injecting Gaussian noise. Then enter a deterministic refinement phase where noise is dropped and an approximate Gauss–Newton correction $(J^{\top}J)^{-1}J^{\top}\nabla_{q}E$ is applied to accelerate convergence.
A.4 Task Definition in Simulations
Task 1: Here, we impose two constraints: the relative pose between the two end-effectors and the position of their midpoint. We define the constraint by specifying the relative transformation matrix $T=H_{1}^{-1}H_{2}$ between the two end-effectors, where in this task $T=\begin{bmatrix}-1&0&0&0\\ 0&1&0&0\\ 0&0&-1&0.7\\ 0&0&0&1\end{bmatrix}$. And the midpoint constraint formed by the two end-effectors is defined by specifying its position $t_{mid}=(x,\,y,\,z)$ in the world coordinate frame.
Task 2: Based on task 1, we added a spherical obstacle to ensure that the robot will not collide with the obstacle. We define obstacles by specifying a sphere’s center position and radius, and during inference we compute the robot’s distance to the sphere to obtain a collision cost, thereby avoiding collisions with the obstacle. Here, we place the obstacle directly above one of the robots, which significantly reduces the robot’s feasible workspace. For the relative pose and midpoint constraints, we use the same settings as in Task 1.
[…]
Task 4: Here given the relative Pose between the two end effectors, and ensure that the midpoint formed by the two end effectors falls within the specified range. We define the valid region for the midpoint by specifying the diagonal corners of a rectangle, i.e. $c_{1}=(x_{1},\,y_{1},\,z_{1})$, $c_{2}=(x_{2},\,y_{2},\,z_{2})$. Here, we assume $z_{1}$ and $z_{2}$ are identical, so that the midpoint is constrained to lie within a horizontal rectangular frame. Similarly, we also impose orientation constraints on the end-effector; here, our constraint is a full rotation constraint, meaning that the rotation angles about each axis are fully specified.
Task 5: Compared to Task 4, we replace the manually specified boundary with input point‐cloud data, so that the sampled points should lie on the surface of the object. As shown in Figure 5, the red dots denote the midpoints formed by the two end-effectors, while the gray dots represent the acquired point-cloud data.
Task 6: Similar to Task 5, but for the end-effector orientation constraint, we only require that it be perpendicular to the object’s surface, which is same as Task4.
A.5 Comparison with Diffusion-CCSP
In our work, we consider Diffusion-CCSP as the primary baseline for comparison. Compared to Diffusion-CCSP, our approach introduces several improvements that lead to stronger performance. While we retain its original network structure, which composes multiple constraint objectives through an energy-based model, for a fair comparison, there are key differences. First, unlike Diffusion-CCSP, our method incorporates the task-specific feature functions from ADCS, which are essential for enhancing generalization. Second, we find that, in particular, Diffusion-CCSP uses fixed composing weights and relying on the ULA sampler during inference, which significantly limit its ability to adapt to different constrained problems and sample effectively. In contrast, our method benefits from transformer-based learnable constraint weighting and a more expressive sequential sampling strategy, resulting in better performance across tasks.
In summary, while CCSP introduced compositionality in diffusion for scene-level object constraints, our work extends this principle to structured, high-dimensional robot constraints. With adaptive weighting, multi-type constraint fusion, and real-world execution, our approach effectively connects abstract constraint composition with practical embodied planning.
A.6 Voxel-based Statistical Method
We also employ a voxel-based statistical method [35] to compute the density of the data distribution. The principle is to uniformly divide the point-cloud’s bounding box into $M^{3}$ small cubes (voxels), count the number $n_{i}$ of points in each voxel, and then assess the dispersion of these counts, as Eq. 8, $\sigma^{2}$ represents the variance, a smaller variance indicates a more uniform data distribution. And $r_{c}$ represent the coverage of the data.
$\sigma^{2}\;=\;\frac{1}{M}\sum_{i=1}^{M}\bigl{(}n_{i}-\bar{n}\bigr{)}^{2},\quad r_{c}=\frac{\sum_{i=1}^{M}\mathbf{1}(n_{i}>0)}{M},\quad\mathbf{1}(n>0)=\begin{cases}1,&n>0,\\ 0,&n\leq 0.\end{cases}$ (8)
A.7 Comparison of Data Distribution from Sampling Results
In this section, we compare the distribution of sampled points obtained by NLP Sampling and our method ADCS under Task 1, 3, 4. For Task 1, 3, we compare the positional distribution of one end-effector. And compare the midpoint distributions between the two end-effectors for Task 4.
From Figure 8, Figure 9, we can observe that in Tasks 1, 3, and 4, the data points sampled by ADCS are more uniformly distributed compared to those obtained by NLP-Sampling.
(a) NLP-Sampling Task 1
(b) ADCS Task 1
(c) NLP-Sampling Task 3
(d) ADCS Task 3
Figure 8: The figure shows the sampling results from NLP-Sampling and ADCS methods for Tasks 1 and 3, illustrating the positional distributions of one end-effector along the $x,y,z$ axes.
(a) NLP-Sampling Task 4
(b) ADCS Task 4
Figure 9: The figure shows the sampling results from NLP-Sampling and ADCS methods for Tasks 4, illustrating the 3D positional distributions of mid points.
A.8 KDE-based Resampling
Under Task 2, we compared KDE-based resampling against uniform-replication resampling, as shown in Figure 10, the KDE-based resampling produces a more uniform distribution along each axis and exhibits fewer voids. Moreover, as shown in Figure 11, the final samples obtained through Gaussian-kernel replication are also more uniformly distributed. We estimate the data distribution density with both voxel size 5 and 10. From Table 6, we can conclude that the KDE-based resampling method achieves a more uniform data distribution both after the resampling stage and in the final sampling results.
(a) Before replication
(b) Uniform replication
(c) KDE-based replication
Figure 10: This figure illustrates the one end-effector’s position distribution densities along the x, y, and z axes before replication, after uniform replication, and after Gaussian-kernel-based replication.
(a) Uniform replication sampling result
(b) KDE-based replication sampling result
Figure 11: This figure shows the one end-effector’s position distribution densities along the x, y, and z axes of the final sampling results after applying uniform replication and after applying Gaussian-kernel–based replication.
Data Distribution Uniformity (Density Variance)
After Replication(5) Final Sampled(5) After Replication(10) Final Sampled(10)
Uniform replication 78.0000 79.6028 5.5500 5.0039
KDE replication 50.4320 57.4560 4.9580 2.8893
Table 6: The table presents the variance of the end-effector’s positional density distribution for the uniform-replication and KDE-based-replication methods, both after the resampling stage and after the final sampling. Smaller variance value indicate a more uniform distribution. We performed out these evaluations using voxel sizes of 5 and 10.
A.9 Sequential Sampling
To evaluate the performance gains brought by sequential sampling, we tested ADCS and ADCS without sequential sampling on six different tasks. As shown in Table 7, for each task under the same number of iterations, sequential sampling consistently yields performance improvements—especially on tasks with smaller feasible regions, such as Task 4 and Task 5, where sequential sampling delivers substantial gains.
Relative Position Error (mm) Mid Point Posistion Error (mm) Relative Rotation Error (rad) End-effector Rotaion Error (rad)
Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean Median Q3
Task 1
ADCS without sequential sampling 7.1670 0.0606 0.2547 5.3705 0.1401 0.4113 0.0090 0.0022 0.0025 - - -
ADCS (ours) 0.4746 0.0191 0.0972 0.2697 0.0561 0.1308 0.0008 0.0005 0.0006 - - -
Task 2
ADCS without sequential sampling 24.0062 0.2334 21.3127 23.3976 0.4774 19.7027 0.0359 0.0025 0.0276 - - -
ADCS (ours) 3.1515 0.0221 0.1004 1.9491 0.0547 0.1294 0.0050 0.0003 0.0004 - - -
Task 3
ADCS without sequential sampling 26.3824 5.0637 27.0838 17.0468 2.4057 15.7216 0.0361 0.0070 0.0393 0.0170 0.0063 0.0175
ADCS (ours) 4.4175 1.7281 4.8418 1.3967 0.6189 1.6780 0.0058 0.0025 0.0067 0.0027 0.0018 0.0034
Task 4
ADCS without sequential sampling 118.5029 63.7327 183.7023 35.5320 16.5970 58.8160 0.1338 0.0473 0.2071 0.1322 0.0700 0.1990
ADCS (ours) 2.1288 0.7755 1.9190 0.5845 0.2435 0.6247 0.0021 0.0009 0.0021 0.0013 0.0007 0.0015
Task 5
ADCS without sequential sampling 72.1891 37.5377 89.3134 83.6005 55.5400 136.1251 0.0773 0.0413 0.0950 0.1223 0.0725 0.1526
ADCS (ours) 7.9545 6.4010 8.2982 0.7715 0.5405 0.7729 0.01375 0.0121 0.0160 0.0304 0.0286 0.0324
Task 6
ADCS without sequential sampling 42.1593 16.4978 54.3672 31.8534 2.7618 27.7955 0.0540 0.0204 0.0658 0.0551 0.0426 0.0642
ADCS (ours) 5.1362 1.7652 4.3743 1.6328 0.5571 1.3955 0.0076 0.0037 0.0072 0.0128 0.0110 0.0154
Table 7: Comparison ADCS and ADCS without sequential sampling under 6 tasks.
A.10 Real-World Experiments
Task 1: A constraint is defined such that the midpoint between the two arms must lie at a fixed position. The sampled points are expected to fall on the same location.
Task 2 and 3: A bounding box is defined such that the sampled midpoints fall within it. We manually define a signed distance function $\mathrm{SDF}(x)$ to enforce this constraint, i.e., ensuring that $\mathrm{SDF}(x)\leq 0$. Specifically, we define a circular bounding box so that the sampled points lie within the circle. Task 2 involves sampling points on a horizontal plane, while Task 3 involves sampling on an inclined surface, meaning the orientation constraints for the end-effectors are different.
Tasks 4 and 5: These are similar to Tasks 2 and 3, except that the circular bounding box is replaced with a rectangular one.
Task 6: Given mesh models of several letters, we convert them into point clouds and constrain the sampled midpoints to lie on these letters. The final result is that the four letters ‘CORL’ are stippled.
Task 7 and 8: Unlike Task 6, here the point cloud information of the object is directly obtained through a camera. The sampled midpoints are constrained to lie on the top surface of the object. Task 7 involves sampling on the top surface of a box, while Task 8 involves sampling on the top surface of a cylinder.
In Table 8, we present the constraint error and the ratio of valid points for each task. An Figure 12 shows the sampling results obtained for each task. Note that each task was tested ten times, and in every task, 500 points were generated through sampling. We consider a point valid if its positional error is less than 3 mm and its orientational error is less than 0.01 rad. As shown, the probability of valid points exceeds 80% in most tasks. Even in tasks with more limited feasible regions, such as Task 1 and Task 6, the valid point probability still exceeds 60%.
Figure 12: Sampling layout of 8 different real-world tasks.
Relative Position Error (mm) Mid Point Posistion Error (mm) Relative Rotation Error (rad) End-effector Rotaion Error (rad) Valid Points Ratio (%)
Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean Median Q3 Mean
Task 1 8.0711 2.2923 6.9647 1.7398 0.9638 1.6872 0.0104 0.0025 0.0074 0.0048 0.0022 0.0042 72.60
Task 2 0.4135 0.0280 0.0706 0.1232 0.0010 0.0024 0.0006 9.06e-05 0.0001 0.0004 2.01e-05 4.27e-05 98.40
Task 3 6.1163 0.0531 0.2818 0.6128 0.0005 0.0018 0.0072 9.10e-05 0.0003 0.0068 3.51e-05 0.0001 88.40
Task 4 2.9277 0.0396 0.1541 0.3672 0.0011 0.0097 0.0025 9.74e-05 0.0002 0.0026 4.21e-05 0.0002 90.60
Task 5 3.8304 0.0630 0.6734 0.5969 0.0009 0.0312 0.0053 0.0001 0.0006 0.0036 4.58e-05 0.0007 84.80
Task 6 3.9650 1.6277 3.8562 0.0574 0.0223 0.0316 0.0065 0.0042 0.0085 0.0080 0.0060 0.0101 66.50
Task 7 3.3118 1.4832 2.0677 0.0489 0.0244 0.0387 0.0052 0.0040 0.0055 0.0055 0.0030 0.0063 82.45
Task 8 3.5339 1.6820 2.3240 0.0469 0.0010 0.0282 0.0059 0.0043 0.0058 0.0068 0.0038 0.0087 80.70
Table 8: The constraint errors and the ratio of valid points of different tasks in the real-world experiment.